# Multi-Step Reasoning with Large Language Models, a Survey

### v2

@@ 3.3.2. Ensemble Strategy
The second kind of reasoning control is based on an ensemble of (sequences of) reasoning steps. For most problems, multiple different options exist for the next step. When all or some of these are generated and evaluated, then the consensus result can be reported as the outcome of an ensemble of steps. Self-consistency (Wang et al., 2023b) and Self-verification (Weng et al., 2023) (in Section 3.2.1) are popular ensemble approaches to evaluate the results of reasoning steps, in which greedy single-path decoding used in Chain-of-thought prompting is replaced by a diverse set of paths. Taking this further, Chain-of-experts uses a mixture-of-experts ensemble for complex combinatorial problems (Xiao et al., 2023b). PAL and MathPrompter also use the ensemble approach. The ensemble approach is popular in reasoning with LLM.
@@ 3.3.3. Reinforcement Learning
We start our discussion of control strategies with depth-first and breadth-first search, then go to beam search, and then to full reinforcement learning. A complex reasoning space can be traversed with a search algorithm. Tree-of-thoughts (Yao et al., 2024) uses breadth-first or depth-first search to dynamically follow different reasoning steps. The evaluation part in Tree-of-thoughts is performed with a prompt to the LLM. Together, the trio of generation, evaluation, and control allow systematic exploration of the space of reasoning steps with look-ahead and backtracking. The authors compare their approach to Chain-of-thought and Self-consistency on the Game of 24, Creative writing, and Mini crossword, achieving an accuracy of 74% on a Game of 24 task. The other tasks are evaluated qualitatively. Figure 5 illustrates the different reasoning structures. [^3] A similarly named approach is Graph-of-thoughts (Besta et al., 2024). Graph-of-thoughts allows more general reasoning graphs, providing a formal framework, where the different elements can then be specified manually.
Figure 5. Reasoning structure of Chain-of-Thought (b), Self-Consistency (c), Tree-of-Thoughts (d) and Buffer of Thoughts (e)
Another approach, Buffer-of-thoughts (Yang et al., 2024), goes a step towards metareasoning (Gao et al., 2024). It introduces a meta-buffer that stores high-level thought-templates. These thought-templates are derived from a variety of tasks. Figure 5 compares the Buffer-of-thoughts approach to other approaches such as Chain-of-thought and Tree-of-thoughts. Buffer-of-thoughts outperforms other methods in puzzles such as Game of 24 (by 11%) and checkmating (by 51%). Thought templates are related to metacognition (thinking about thinking), which is further discussed in Section 4.3.3.
@@
Reinforcement learning is another step in the sophistication of optimization algorithms. It learns by interactive sampling, improving its policy based on rewards from the environment (Sutton and Barto, 2018). To use reinforcement learning, the reasoning problem is formulated as a Markov Decision Process: the agent-algorithm creates a prompt (an action), to sample a step ($t$) and get an answer (state, reward) from the environment-model. The answer can then be used to improve the prompt (next action), using the rewards to improve its policy of best actions for each state. The approaches that use reinforcement learning also do so in the form of an external algorithm.
@@
Another approach that combines reinforcement learning and LLMs is ReAct (Yao et al., 2023). Most works focus on reasoning by the LLM, not on actions by an agent. The goal of ReAct is to combine progress in reasoning with action plan generation. (Or, to put it differently, other approaches use RL to improve LLM-reasoning, ReAct uses LLMs to improve RL agent policies.) ReAct uses Chain-of-thought prompt-learning as part of an RL framework that also uses external knowledge sources (Wikipedia) and finetuning; for error reduction, grounding, and for reducing hallucination. The framework allows hand-written prompts. On two interactive decision making benchmarks (Alfworld and WebShop), ReAct outperforms imitation and reinforcement learning methods by an absolute success rate of 34% and 10% respectively.
@@
The ReAct work has been developed further. Reflexion (Shinn et al., 2024) creates AI agents that learn by reflecting on failures and enhance their results, much like humans do. Reflexion uses three language models: actor, evaluator, and reflector. It works as follows: (1) an actor generates text and actions, (2) an evaluator model scores the outputs produced by the actor, and (3) a self-reflection model generates verbal reinforcement cues to assist the actor to self-improve (see Figure 6). For the actor, Chain-of-thought and ReAct can be used. Reflexion is evaluated on decision-making, reasoning, and coding tasks. Improvements of 10-20 percentage points are reported.
@@
Figure 6. Architecture of Reflexion (Shinn et al., 2024), showing a close resemblance to the agent/environment structure of reinforcement learning
@@
The reinforcement learning approaches that we discussed so far—React, Self-refine, Tree-of-thoughts, Buffer-of-thoughts, Reflexion—use external algorithms to manage state for the prompt improvement loop. A more elegant solution would be to perform reinforcement learning fully in-context, within-prompt. Indeed, Krishnamurthy et al. (2024) explicitly asked the question whether LLMs can explore in context. This is the goal of the Algorithm-of-thoughts (AoT) approach (Sel et al., 2023). Following work by Lee et al. (2023); Wang et al. (2024c); Demircan et al. (2024) that showed that transformer architectures can be pretrained and finetuned to perform in-context reinforcement learning, AoT aims to do so in general LLMs such as GPT-4, Claude and Gemini 1.5. They achieve results comparable to Tree-of thoughts on GSM8K, StrategyQA, and Crosswords. Achieving these results with in-context RL is a promising result for in-context learning. Other search-like in-context algorithms are studied by Schultz et al. (2024) and Kempinski et al. (2025) (Game-of-thoughts). Further work in games finds that LLMs struggle with the knowing-doing gap and the difference between using and mentioning game concepts (Paglieri et al., 2024; van Wetten et al., 2025; Cloos et al., 2024).
@@
To conclude this overview of reinforcement learning approaches, we discuss an application in the games domain. Voyager (Wang et al., 2024a) is an agent for the game of Minecraft that uses an iterative prompting mechanism that generates code for embodied control. The agent includes self-verifcation an a skill library to maximize exploration. The goal of is to discover diverse items in Minecraft, a form of novelty search (Eysenbach et al., 2019). Voyager performs well, reaching high scores by acquiring many tools (see Figure 7) 15 times faster than the baseline.
@@
The applications of reinforcement learning in LLM reasoning are many, and the connections run deep (Pternea et al., 2024). Wang et al. (2024b) use the similarity between RL timesteps and LLM reasoning steps to jointly train a value function together with the LLM policy by optimizing the soft Bellman equation, achieving 85% accuracy on GSM8K and 81% on Alfworld. Guo et al. (2025); Du et al. (2025) replace supervised finetuning by reinforcement learning, integrating it with reasoning, achieving large efficiency gains in the training pipeline in the DeepSeek r1 and Kimi models.
@@ 4. Discussion
We have reviewed many reasoning approaches. It is now time to reflect on the approaches, to discuss limitations, and to look for promising areas of future work. First we discuss how to choose an approach for an application. Next, we discuss issues concerning hallucination, faithful reasoning, and scaling. Then we discuss what LLMs can and cannot do. Then, we highlight connections with sequential decision processes, metacognition, and programming languages, and we end with a research agenda.
@@ 4.1. Matching Methods and Applications
This survey has discussed different approaches, applications, and benchmarks. It is often difficult to directly compare benchmark results between individual papers, due to differences in measurement setup. For example, Kamoi et al. (2024) critically compare self-reflection studies. We can provide, however, general guidance on what approach to use for different types of applications.
For simple reasoning situations, that require a single sequence of reasoning steps, Chain-of-thought is typically used. When the error rate is high, verification methods such as Self-consistency are popular. Indeed, this combination is used in many modern reasoning LLMs.
For combinatorial puzzles and games that would traditionally require backtracking, tree-search or reinforcement learning, solution methods need to be able to keep track of the enumeration state. However, an LLM without an external algorithm often fails to keep track of this state correctly between calls to the model. An approach such as Tree-of-thoughts, or ReAct, should be considered. Here, the state is managed externally, and the LLM is called for application-dependent functional processing. Various in-context reinforcement learning methods are being developed where state is managed by the LLM, as in, for example, Algorithm-of-thoughts. When performance is not sufficient, finetuning or pretraining for the domain at hand can be used.
For complex problems, where a sequence of steps or external control fail to perform well, other reasoning approaches are needed. We note that LLMs are good at generating computer code for many known problems (Chen et al., 2021). Hence, when the preceding approaches do not work, the LLM can be asked to generate a program for the problem. In this approach, the LLM acts as a coding agent for its user. The LLM generates a program or a sequence of commands, which are then executed by an external tool, such as an interpreter, a robot, a planner, or a logic solver. In robotics, vision-language-action models have achieved impressive results (Kim et al., 2024), but also strong results in logistics (Bohnet et al., 2024), finance and medicine are reported (Plaat et al., 2025).
@@ 4.2.1. Faithfulness
Chain-of-thought approaches prompt a language model to take certain steps to solve the problem that the prompt specifies. One can ask the question, whether those steps are indeed the steps that the model has followed (faithful reasoning) or whether it took another road to arrive at the same answer (unfaithful reasoning). A few studies measure the faithfulness of reasoning with LLMs. Lanham et al. (2023) notes that just like organic reasoners, a model’s reasoning may be post-hoc, it may be constructed after a certain conclusion has been found. By deliberately adding mistakes to the chain of thought, the authors measure the faithfulness of the model. They find a wide variation of post-hoc reasoning, with a tendency of larger models to be less faithful. Like regular LLMs, when not properly grounded, (Chain-of-thought) reasoning suffers from hallucination (Huang et al., 2025).
Another study adds deliberate bias to the prompt. For example, in a multiple-choice setting, they always make answer (A) the correct answer (Turpin et al., 2024). They find that a bias towards wrong answers can cause significant drops in accuracy, and that models frequently generate Chain-of-though explanations rationalizing wrong answers. To address issues of faithfulness, Lyu et al. (2023); Xu et al. (2024a) propose Faithful-chain-of-thought. This approach involves two stages. First, the natural language query is translated into a formal symbolic language. Second, the problem-solving stage processes the formal language, and can explain the reasoning steps it has thus taken. For the symbolic language, Python, Datalog, or PDDL is suggested. Another approach, mechanistic interpretability, studies methods to target individual representations inside the LLM, to see if the expected behavior occurs in practice (Rai et al., 2024; Bereska and Gavves, 2024; Chen, 2025).
@@ 4.3.1. What Can LLMs Do?
A taxonomy of generate-evaluate-control is able to describe the structure of the current LLM reasoning literature. Furthermore, the accuracy of the reasoning chains can be improved with ensemble methods, and self-verification. Hallucination can be reduced by grounding the model with external models, such as for robotic affordances, and information retrieval from search engines and Wikipedia. Going a step further, using external control algorithms as scaffolding (such as search or reinforcement learning), dynamic prompts can use the LLMs to perform complex and dynamic reasoning patterns. Note that the reasoning control is now two layers away from the core LLM: an external control algorithm, on top of in-context-learning, dynamically generating prompts for the LLM. This is reasoning with prompts with LLMs, not by.
At this point, it is interesting to note the confluence of the two schools of classical artificial intelligence, symbolic and connectionist. [^4] Reasoning and planning have been studied since the start of artificial intelligence, starting with logic and reasoning (Newell and Simon, 1961), search algorithms in puzzles and board games (Korf, 1999; Plaat, 2020), robot planning (Fikes and Nilsson, 1971), classical machine learning such as decision trees and support vector machines (Flach, 2012; Breiman, 2001; Cortes and Vapnik, 1995), through knowledge representation and the semantic web (Van Harmelen et al., 2008). Ever since the success of the connectionist approach (LeCun et al., 2015; Goodfellow et al., 2016) (deep learning, including LLMs) researchers have tried to join the two approaches. Search and reinforcement learning are rooted in the symbolic tradition, while LLMs are rooted in the connectionist tradition. The literature in this survey combines the two traditions. High performance reasoning is created with a (symbolic) searcher/learner on top of a (connectionist) LLM. In other fields similar combinations can be seen (for example, AlphaFold (Bryant et al., 2022; Jumper et al., 2021) and retrosynthesis of molecules (Segler et al., 2018)). The LLM helps ground symbolic reasoning methods in language; symbolic methods help create prompts that let the LLM perform dynamic reasoning.

### v3

@@ 3.3.2. Ensemble Strategy
The second kind of reasoning control is based on an ensemble of (sequences of) reasoning steps. For most problems, multiple different options exist for the next step. When all or some of these are generated and evaluated, then the consensus result can be reported as the outcome of an ensemble of steps. Self-consistency (Wang et al., 2023c) and Self-verification (Weng et al., 2023) (in Section 3.2.1) are popular ensemble approaches to evaluate the results of reasoning steps, in which greedy single-path decoding used in Chain-of-thought prompting is replaced by a diverse set of paths. Taking this further, Chain-of-experts uses a mixture-of-experts ensemble for complex combinatorial problems (Xiao et al., 2023b). Program-aided-language and MathPrompter also use the ensemble approach. The ensemble approach is popular in reasoning with LLM.
@@ 3.3.3. Reinforcement Learning
We start our discussion of control strategies with depth-first and breadth-first search, then go to beam search, and then to full reinforcement learning. A complex reasoning space can be traversed with a search algorithm. Tree-of-thoughts (Yao et al., 2024) uses breadth-first or depth-first search to dynamically follow different reasoning steps. Tree-of-thoughts calls the LLM with two different types of prompt: to generate the sub-problems, and to evaluate them. Together, the trio of generation (LLM-prompt-1), evaluation (LLM-prompt-2), and control (search-algorithm) allow systematic exploration of the space of reasoning steps with look-ahead and backtracking. The authors compare their approach to Chain-of-thought and Self-consistency on the Game of 24, Creative writing, and Mini crossword, achieving an accuracy of 74% on a Game of 24 task. The other tasks are evaluated qualitatively. Figure 5 illustrates the different reasoning structures. [^4] A similarly named approach is Graph-of-thoughts (Besta et al., 2024). Graph-of-thoughts allows more general reasoning graphs, providing a formal framework, where the different elements can then be specified manually.
Figure 5. Reasoning structure of Simple prompting (a), Chain-of-thought (b), Self-Consistency (c), Tree-of-Thoughts (d) and Buffer of Thoughts (e); Prompts in (a) consist of an instruction, and, perhaps, few-shot examples of answers, providing no reasoning guidance; Chain-of-thought (b) guides the model with few-shot examples of reasoning steps; Self-Consistency (c) takes the majority vote of existing answers in the model; Tree-of-thoughts (d) uses an external prompt-optimizing algorithm to guide the model to perform a tree search over reasoning steps, exploring multiple alternatives; Buffer-of-thoughts (e) uses a problem distiller that stores high-level thoughts distilled from different problems in a meta-buffer, which is actively managed for capacity
Another approach, Buffer-of-thoughts (Yang et al., 2024), goes a step towards metareasoning (Gao et al., 2024). It introduces a meta-buffer that stores high-level thought-templates. These thought-templates are derived from a variety of tasks by a problem distiller. They are then instantiated for specific tasks. To address size restrictions of the LLM context window, the meta-buffer is managed for capacity and thoughts can be combined. Figure 5 compares the Buffer-of-thoughts approach to other approaches such as Chain-of-thought and Tree-of-thoughts. Buffer-of-thoughts outperforms other methods in puzzles such as Game of 24 (by 11%) and checkmating (by 51%). Thought templates are related to metacognition (thinking about thinking), which is further discussed in Section 4.3.3.
@@
Reinforcement learning is another step in the sophistication of optimization algorithms. It learns by interactive sampling, improving its policy based on rewards from the environment (Sutton and Barto, 2018). To use reinforcement learning, the reasoning problem is formulated as a Markov Decision Process: the agent-algorithm creates a prompt for a next step (an action), to sample a step ($t$) and get an answer (state, reward) from the environment-model. The answer can then be used to improve the prompt (next action), using the rewards to improve its policy of best actions for each state. The approaches that use reinforcement learning also do so in the form of an external algorithm.
@@
Another approach that combines reinforcement learning and LLMs is ReAct (Yao et al., 2023). However, ReAct does so in a different way. Most works focus on reasoning by the LLM, not on actions by an agent. The goal of ReAct is to combine progress in reasoning with action plan generation. (Or, to put it differently, other approaches use RL to improve LLM-reasoning, ReAct uses LLMs to improve RL agent policies.) ReAct uses Chain-of-thought prompt-learning as part of an RL framework that also uses external knowledge sources (Wikipedia) and finetuning; for error reduction, grounding, and for reducing hallucination. The framework allows hand-written prompts. On two interactive decision making benchmarks (Alfworld and WebShop), ReAct outperforms imitation and reinforcement learning methods by an absolute success rate of 34% and 10% respectively.
@@
The ReAct work has been developed further. Reflexion (Shinn et al., 2024) creates AI agents that learn by reflecting on failure and enhance their results, much like humans do. Reflexion uses three language models: actor, evaluator, and reflector. It works as follows: (1) an actor generates text and actions, (2) an evaluator model scores the outputs produced by the actor, and (3) a self-reflection model generates verbal reinforcement cues to assist the actor to self-improve (see Figure 6). For the actor, Chain-of-thought and ReAct can be used. Reflexion is evaluated on decision-making, reasoning, and coding tasks. Improvements of 10-20 percentage points are reported.
@@
Figure 6. Architecture of Reflexion (Shinn et al., 2024), showing a close resemblance to the agent/environment structure of reinforcement learning (Sutton and Barto, 2018)
@@
The reinforcement learning approaches that we discussed so far—React, Self-refine, Tree-of-thoughts, Buffer-of-thoughts, Reflexion—use external algorithms to manage state for the prompt improvement loop. A more elegant solution would be to perform reinforcement learning fully in-context, within-prompt. Indeed, Krishnamurthy et al. (2024) explicitly asked the question whether LLMs can explore in-context. This is the goal of the Algorithm-of-thoughts (AoT) approach (Sel et al., 2023). Following work by Lee et al. (2023); Wang et al. (2024c); Demircan et al. (2024) that showed that transformer architectures can be pretrained and finetuned to perform in-context reinforcement learning, AoT aims to do so in general LLMs such as GPT-4, Claude and Gemini 1.5. They achieve results comparable to Tree-of thoughts on GSM8K, StrategyQA, and Crosswords. Achieving these results with in-context RL is a promising result for in-context learning. Other search-like in-context algorithms are studied by Schultz et al. (2024) and Kempinski et al. (2025). Further works in games find that LLMs struggle with the difference between generating an algorithm for a problem, and executing that algorithm correctly (the knowing-doing gap (Paglieri et al., 2024; Ruoss et al., 2024)), and struggle with the difference between using and mentioning game concepts (van Wetten et al., 2025; Cloos et al., 2024). Work on implicit reasoning is ongoing, Li et al. (2025) provide a survey.
@@
To conclude this overview of reinforcement learning approaches, we discuss an application in the games domain. Voyager (Wang et al., 2024a) is an agent for the game of Minecraft that uses an iterative prompting mechanism that generates code for embodied control. The agent includes self-verifcation and a skill library to maximize exploration. The goal is to discover diverse items in Minecraft, a form of novelty search (Eysenbach et al., 2019). Voyager performs well, reaching high scores by acquiring many tools (see Figure 7) 15 times faster than the baseline.
@@
The applications of reinforcement learning in LLM reasoning are many, and the connections run deep (Pternea et al., 2024). Wang et al. (2024b) use the similarity between RL timesteps and LLM reasoning steps to jointly train a value function together with the LLM policy by optimizing the Bellman equation, achieving 85% accuracy on GSM8K and 81% on Alfworld. Guo et al. (2025); Du et al. (2025) replace supervised finetuning by reinforcement learning, integrating it with reasoning, achieving large efficiency gains in the training pipeline in the DeepSeek r1 and Kimi models.
@@ 4. Discussion
We have reviewed many reasoning approaches. It is now time to reflect, to discuss limitations, and to look for promising areas of future work. First we discuss how to choose an approach for an application. Next, we discuss issues concerning hallucination, faithful reasoning, and scaling. Then we discuss what LLMs can and cannot do. We highlight connections with sequential decision processes, metacognition, and programming languages, and we end with a research agenda.
@@ 4.1. Matching Methods and Applications
This survey has discussed different approaches, applications, and benchmarks. It is often difficult to directly compare benchmark results between individual papers, due to differences in measurement setup. For example, Kamoi et al. (2024) critically compare self-reflection studies. Despite the challenges in comparing benchmark results, we can provide general guidance on what approach to use for different types of applications.
For simple reasoning situations, that require a single sequence of reasoning steps, Chain-of-thought is typically used. When the error rate becomes too high, verification methods such as Self-consistency are popular. Indeed, this combination is used in many modern reasoning LLMs such as DeepSeek and GPT-o1 (Guo et al., 2025; Du et al., 2025; Lambert et al., 2024; Huang et al., 2024).
For combinatorial puzzles and games that would traditionally require backtracking, tree-search or reinforcement learning solution methods must keep track of the enumeration state. However, an LLM without an external algorithm often fails to keep track of this state correctly between calls to the model. An approach such as Tree-of-thoughts, or ReAct, should be considered. Here, the state is managed externally, and the LLM is called for application-dependent functional processing. Various in-context reinforcement learning methods are being developed where state is managed in-context, as in, for example, Algorithm-of-thoughts. When performance is not sufficient, finetuning or pretraining for the domain at hand can be used.
For complex problems, where a sequence of steps, or external control, fail to perform well, other reasoning approaches are needed. The LLM can be asked to generate a program for the problem. In this approach, the LLM acts as a coding agent for its user. The LLM generates a program or a sequence of commands, which are then executed by an external tool, such as an interpreter, a robot, a planner, or a logic solver. In robotics, vision-language-action models have achieved impressive results (Kim et al., 2024), but also strong results in logistics (Bohnet et al., 2024), finance and medicine are reported (Plaat et al., 2025).
@@ 4.2.1. Faithfulness
Chain-of-thought approaches prompt a language model to take certain steps to solve the problem that the prompt specifies. One can ask the question whether those steps are indeed the steps that the model has followed (faithful reasoning) or whether it took another road to arrive at the same answer (unfaithful reasoning). A few studies measure the faithfulness of reasoning with LLMs. Lanham et al. (2023) notes that just like organic reasoners, a model’s reasoning may be post-hoc, it may be constructed after a certain conclusion has been found. By deliberately adding mistakes to the chain of thought, the authors measure the faithfulness of the model. They find a wide variation of post-hoc reasoning, with a tendency of larger models to be less faithful. Like regular LLMs, when not properly grounded, (Chain-of-thought) reasoning suffers from hallucination (Huang et al., 2025).
Another study adds deliberate bias to the prompt. For example, in a multiple-choice setting, they always make answer (A) the correct answer (Turpin et al., 2024). They find that a bias towards wrong answers can cause significant drops in accuracy, and that models frequently generate Chain-of-though explanations rationalizing wrong answers. To address issues of faithfulness, Lyu et al. (2023); Xu et al. (2024a) propose Faithful-chain-of-thought. This approach involves two stages. First, the natural language query is translated into a formal symbolic language. Second, the problem-solving stage processes the formal language, and can explain the reasoning steps it has thus taken. For the symbolic language, Python, Datalog, or PDDL is suggested. Another approach, mechanistic interpretability, studies methods to target individual representations inside the LLM, to see if the expected behavior occurs in practice (Rai et al., 2024; Bereska and Gavves, 2024; Chen et al., 2025).
@@ 4.3.1. What Can LLMs Do?
A taxonomy of generate-evaluate-control is able to describe the structure of the current LLM reasoning literature. Furthermore, the accuracy of the reasoning chains can be improved with ensemble methods, and self-verification. Hallucination can be reduced by grounding the model with external models, such as for robotic affordances, and information retrieval from search engines and Wikipedia. Going a step further, using external control algorithms as scaffolding (such as search or reinforcement learning), dynamic prompts can use the LLMs to perform complex and dynamic reasoning patterns. Note that the reasoning control is now outside the core LLM: an external control algorithm, on top of in-context-learning, dynamically generating prompts for the LLM.
At this point, it is interesting to note the confluence of two schools of thought in artificial intelligence: symbolic and connectionist. [^5] Reasoning and planning have been studied since the start of artificial intelligence, starting with logic and reasoning (Newell and Simon, 1961), search algorithms in puzzles and board games (Korf, 1999; Plaat, 2020, 1996), robot planning (Fikes and Nilsson, 1971), classical machine learning such as decision trees and support vector machines (Flach, 2012; Breiman, 2001; Cortes and Vapnik, 1995), through knowledge representation and the semantic web (Van Harmelen et al., 2008). Ever since the success of the connectionist approach (Brooks, 1990; LeCun et al., 2015; Goodfellow et al., 2016) (deep learning, including LLMs) researchers have tried to join the two approaches (Tesauro et al., 1995; Silver et al., 2017; Wang et al., 2023b). Search and reinforcement learning are rooted in the symbolic tradition, while LLMs and deep learning are rooted in the connectionist tradition. The literature in this survey combines the two traditions. Higher performance reasoning is created with a (symbolic) searcher/learner on top of a (connectionist) LLM. In other fields similar combinations can be seen (for example, AlphaFold (Bryant et al., 2022; Jumper et al., 2021) and retrosynthesis of molecules (Segler et al., 2018)). The LLM helps ground symbolic reasoning methods in language; symbolic methods help create prompts that let the LLM perform dynamic reasoning.